# Obfuscation Based Privacy Preserving Representations are Recoverable Using Neighborhood Information

### v1

@@ 6 Experimental Evaluation
Superpoint[17] SIFT[40]
Geometric Perceptual Geometric Perceptual
Lines CP Lines CP Lines CP Lines CP
Oracle In 0.75 61.4 91.2 54.4 83.4 0.46 0.45 77.6 96.1 72.7 91.8 0.62 0.58
Oracle In. 0.5 63 92.9 45.6 70.3 0.46 0.42 78.8 96.9 63.9 80.2 0.62 0.58
Oracle In. 0.3 57.8 87.8 23.7 38.4 0.45 0.4 73.1 92.0 32.3 43.6 0.61 0.53
Oracle In 0.2 45.1 70.8 13.3 23.1 0.40 0.4 56.9 74.8 15.9 24.6 0.57 0.53
Oracle In. 0.1 22.9 39.0 6.7 13.3 0.33 0.4 26.2 40.0 6.9 13.2 0.51 0.52
Table 3: End-to-end attack evaluation. Geometric and perceptual evaluation of the recovery when using neighbors estimated by our network described in Sec.5 (last row) for two different types of keypoint detectors and extractors - Superpoint[17] and SIFT[40]. The performance when using oracle provided neighborhoods of different qualities is provided for comparison. The recovery of neighborhoods is observed to be much more effective using Superpoint [17] descriptors as compared to SIFT[40].
@@
Once the approximate positions of the obfuscated points are recovered, they are fed together with their descriptors to an inversion network [22, 49] to generate images of the scene. This assumes that the geometric obfuscation preserves the descriptor, which is a valid assumption in the context of visual localization. This is true for all the obfuscations discussed in this paper except for PPL and its variant PPL+ [36]. These obfuscations map a pair of points and the corresponding descriptors to the same line without preserving the mapping between the points and their descriptors. We again use the neighborhood information to recover the original mapping between a point and a descriptor (see supp. mat. for details).
Datasets and metrics.  We evaluate on the two indoor datasets 7-scenes [61] and 12-scenes [69], and the outdoor dataset Cambridge [31]. Results on 12-scenes are included in the supp. mat as they follow the same trend as results on 7-scenes. Similarly, results for SuperPoint [17] are left for supp. mat as they follow a similar trend as the results using SIFT [40]. We report the geometric accuracy as the fraction of points recovered within chosen error thresholds of 5, 10, and 25 pixels in 2D, and 10, 25, and 50 cm in 3D depending on whether the scene is indoor (small-scale) or outdoor (large-scale). We compare the quality of the images generated from the recovered points against the ones generated from the original points by comparing their respective similarity to the real image. The similarity is computed with standard perceptual metrics: SSIM, PSNR and LPIPS [75]. We report the last two metrics in the supp. material.
Geometric evaluation.  We report the geometric accuracy for 2D and 3D in Table 1 and Table 2, respectively. The proposed generic recovery method can consistently recover the points within a few pixels in the 2D case and within a few cms in 3D maps. Note that for 2D and 3D line-based obfuscations the performance can peak when the neighborhood is not perfect, i.e., when the inlier ratio is lower than 1.0, although the variation is small. This is because our robust method identifies outliers in the neighborhood and filters them out. Although outlier-free, such a neighborhood has a different size and different elements compared to the perfect neighborhood with inlier ratio 1.0 so the optimization can estimate slightly different recovered points.
Perceptual evaluation.  All perceptual metrics show consistent results so we report SSIM only in Tab. 4. ‘GT’ is the baseline SSIM between the real image and the one generated from the original points, which can be interpreted as an upper bound for the SSIM between the real image and the one generated from the recovered points. The difference from this bound is higher for 2D than for 3D. This is because even a high geometric error in 3D can reduce to very few pixels upon projection while the recovery from obfuscations in 2D leads to several pixels of error. This larger error in the sparse set of keypoint locations, which is an input to the inversion network, leads to a larger gap in perceptual metrics.
@@
An interesting observation is that the network learns the neighborhood more easily for SuperPoint [17] than for SIFT [40] as indicated by the accuracy gap between the two: in the first case, the network leads to neighborhoods with accuracies between 70% and 80% for $K\in[10,100]$ while for SIFT, the accuracy remains around 35%. Thus one could argue that SIFT is more privacy-preserving than SuperPoint, although there is no guarantee that better neighborhood estimators for SIFT will not become available in the future. Moreover, SIFT typically achieves lower localization performance than SuperPoint [55] and sacrificing performance for privacy might not be a satisfying solution in all scenarios.
Even though the images are not perfectly inverted from the recovered points, the outline and the objects in the scene are recognizable. These results highlight an important limitation of pure geometric obfuscations and support two claims made in the paper: i) the neighborhood information can be learned from the descriptors preserved by the obfuscations and reiterates that geometric obfuscations alone are not as privacy-preserving as they claim. One needs to also prevent neighborhood information from being inferred from the obfuscations. ii) it shows that the proposed proof of concept to compute the neighborhood information is already sufficient for the proposed point recovery to be applicable. We expect that more complex neighborhood learning will lead to better results. This calls for further work on obfuscating the descriptors to prevent learning neighborhoods or other private content, and fusing geometric and descriptors obfuscations.
Discussion.  The results reveal that the proposed recovery method performs well even if the neighborhoods contain significant fractions of outliers. Fig. 2 and Fig. 3 further show that images generated from recovered points can reveal potentially private user content, which is particularly true for line-based obfuscations [63, 64, 36]. The geometric constraints of parallel planes [26] make recovery difficult, but neighborhoods with reasonable inlier ratios make the plane obfuscation also susceptible to the proposed recovery. The same holds for Coordinate Permutation [45]: neighborhoods with inlier ratios of 0.5 or more are enough to enable recovery accurate enough to reveal identifiable scene content. The network described in Sec. 5 can produce such informative neighborhoods even with its simple design. We expect methods in future works to improve the estimation of neighborhoods from descriptors, further highlighting the discussed vulnerability of obfuscation schemes.
@@ 7 Conclusion
In this work, we highlight a common vulnerability of all geometry-based obfuscation techniques that have so far been presented as privacy-preserving representations. We present a simple optimization-based method that uses knowledge of point neighborhoods to recover point positions from the discussed obfuscation schemes. We show the robustness of our method and analyze the recovery accuracy by using oracle-provided neighborhoods with varying inlier ratios. Finally, using a neural network that learns to identify local feature descriptors co-occurring across scenes, we show that it is possible to estimate these neighborhoods from the descriptors associated with points. The inverted images from the recovered point positions reveal private scene content, highlighting the drawback of current methods and the need for guarantees on under which circumstances a method for privacy-preserving localization is indeed privacy-preserving.
Acknowledgements
This work was supported by the Czech Science Foundation (GACR) JUNIOR STAR Grant No. 22-23183M, the Czech Science Foundation (GACR) EXPRO (grant no. 23-07973X), the Chalmers AI Research Center (CHAIR), WASP and SSF. The compute and storage were partially supported by NAISS projects numbered NAISS-2024/22-637 and NAISS-2024/22-237 respectively.
@@ Acknowledgements
Appendix A Appendix
The supplementary material is organized as follows. Sec. A.1 details the point recovery from coordinate permutations [45] and how we estimate which of the coordinates is swapped to transform the recovery from coordinate permutations into a recovery from lines. Sec. A.2 recalls the descriptor ambiguity in paired-point lines obfuscations [36] and how neighborhood information is used to assign descriptors to their original points. Sec. A.3 reports results on the indoor 12-scenes [69] dataset as announced in Section 6. These results are consistent with the ones on the indoor 7-scenes [61] dataset. We also report the geometric and perceptual evaluation for all 3D obfuscations, including the random line obfuscation OLC [63] and the PPL+ variant of the pair-point lifting [36], that are left out of the main paper for the sake of brevity. Sec. A.4 provides additional implementation details related to the nearest-neighbor learning and the image inversion from 2D points.
A.1 Coordinate permutation - Predicting swapped coordinate
@@
A.2 Descriptor Assignment for Paired-Point-Lines
@@ A.4 Implementation Details
For 3D lines [63, 36, 43] and 3D lines made from 3D permutation [45], the 3D points to recover are initialized as the projection of a 3D point "anchor" onto the lines. The 3D anchor is defined as follows: the 3D lines are projected onto a plane. We sample a set of intersections between the resulting 2D lines and compute their 2D centroid, which we use as the anchor. In the paper, we use the plane $z=0$ so the centroid has the form $(x,y,0)$ and randomly sample 10K intersection points. Note that the choice for the plane $z=0$ does not necessarily correspond to the ground-plane as the coordinate frames of the scenes are chosen arbitrarily by the authors of the datasets.
@@
For planes, we also project a 3D "anchor" point onto each plane but the anchor point is built differently: it is defined as the 3D point which coordinates are the average of the planes’ offsets associated with that axis, i.e., the $x$ coordinate is the average of offsets $c$ of all planes of the form $x=c$.
@@
[…]

### v2

@@ 6 Experimental Evaluation
Superpoint [16] SIFT [39]
Geometric Perceptual Geometric Perceptual
Lines CP Lines CP Lines CP Lines CP
Oracle In. 0.75 61.4 91.2 54.4 83.4 0.46 0.45 77.6 96.1 72.7 91.8 0.62 0.58
Oracle In. 0.5 63.0 92.9 45.6 70.3 0.46 0.42 78.8 96.9 63.9 80.2 0.62 0.58
Oracle In. 0.3 57.8 87.8 23.7 38.4 0.45 0.40 73.1 92.0 32.3 43.6 0.61 0.53
Oracle In. 0.2 45.1 70.8 13.3 23.1 0.40 0.40 56.9 74.8 15.9 24.6 0.57 0.53
Oracle In. 0.1 22.9 39.0 6.7 13.3 0.33 0.40 26.2 40.0 6.9 13.2 0.51 0.52
Table 4: End-to-end attack evaluation. Geometric and perceptual evaluation of the recovery when using neighbors estimated by our network described in Sec. 5 (last row) for two different types of keypoint detectors and extractors—Superpoint [16] and SIFT [39]. The performance when using oracle-provided neighborhoods of different qualities is provided for comparison. The recovery of neighborhoods is observed to be much more effective using Superpoint [16] descriptors compared to SIFT [39].
@@
The recovered positions of the obfuscated points, together with the descriptors are fed to an inversion network [21, 48] to generate images of the scene. As a valid assumption in the context of visual localization, the descriptor is assumed to be not modified. This is true for all the obfuscations discussed in this paper except for PPL/PPL+ [35]. These obfuscations map a pair of points and the corresponding descriptors to the same line, without preserving the mapping between the points and their descriptors. We again use the neighborhood information to recover this point-descriptor mapping (see supp. mat. for details).
Datasets and metrics.  We evaluate on the two indoor datasets 7-scenes [60] and 12-scenes [68], and the outdoor dataset Cambridge [30]. Results on 12-scenes are included in the supp. mat as they follow the same trend as results on 7-scenes. Similarly, results for SuperPoint [16] are left for supp. mat as they follow a similar trend as the results using SIFT [39]. We report the geometric accuracy as the fraction of points recovered within chosen error thresholds. The threshold is in pixels for 2D obfuscations and in cm for 3D. Larger thresholds are used for larger (outdoor) scenes in 3D. We compare the quality of the images generated from the recovered points against the ones generated from the original points by comparing their respective similarities to the real image. The similarity is computed with standard perceptual metrics: SSIM, PSNR and LPIPS [76]. We report the last two metrics in the supp. material.
Geometric evaluation.  The geometric accuracies for 2D and 3D are reported in Tables 1 and 2, respectively. The proposed generic recovery method can consistently recover the points within a few pixels in the 2D case and within a few cms in 3D maps. Note that for 2D and 3D line-based obfuscations the performance can peak when the neighborhood is not perfect, i.e., when the inlier ratio is lower than 1.0, although the variation is small. This is because our robust method identifies outliers in the neighborhood and filters them out and using fewer close points can result in better accuracy using our method.
Perceptual evaluation.  All perceptual metrics show consistent results so we report only the SSIM in Tab. 3. ‘GT’ is the baseline SSIM between the real image and the one generated from the original points, which can be interpreted as an upper bound for the SSIM between the real image and the one generated from the recovered points. The difference from this bound is higher for 2D than for 3D. This is because even a high geometric error in 3D can reduce to very few pixels upon projection while the recovery from obfuscations in 2D leads to several pixels of error. The larger error in 2D keypoint position estimation leads to worse image reconstructions in case of 2D obfuscations.
@@
An interesting observation is that the network learns the neighborhood more easily for SuperPoint [16] than for SIFT [39] as indicated by the accuracy gap between the two: for SuperPoint [16], the network leads to neighborhoods with acc. between 70% and 80% for $K\in[10,100]$ while for SIFT [39], the accuracy remains around 35%. Thus one could argue that SIFT is more privacy-preserving than SuperPoint, although there is no guarantee that better neighborhood estimators for SIFT will not become available in the future. Moreover, SIFT [39] typically achieves lower localization performance than SuperPoint [54] and sacrificing performance for privacy might not be a satisfying solution in all scenarios.
Even though the images inverted from the recovered points are not perfect, the outline and the objects in the scene are recognizable. These results highlight an important limitation of pure geometric obfuscations and support the two claims made in the paper: i) the neighborhood information can be learned from the descriptors, and reiterates that geometric obfuscations alone are not as privacy-preserving as they claim. One needs to also prevent neighborhood information from being inferred from the obfuscations. ii) it shows that the proposed proof of concept to compute the neighborhood information is already sufficient for the proposed point recovery to be applicable. We expect that more complex neighborhood learning will lead to better results. This calls for potential future work on fusing geometric and descriptor obfuscation to prevent neighborhood recovery.
Discussion.  The results reveal that the proposed recovery method performs well even if the neighborhoods contain significant fractions of outliers. Fig. 2 and Fig. 3 further show that images generated from recovered points can reveal potentially private user content, which is particularly true for line-based obfuscations [62, 63, 35]. The geometric constraints of parallel planes [25] make recovery difficult, but neighborhoods with reasonable inlier ratios make the plane obfuscation also susceptible to the proposed recovery. The same holds for Coordinate Permutation [44]: the additional step of estimating which coordinate was permuted brings in more noise into our method for recovering points. However, neighborhoods with inlier ratios of 0.5 or more are enough to enable recovery accurate enough to reveal identifiable scene content. The network described in Sec. 5 can produce such informative neighborhoods even with its simple design. We expect methods in future works to improve the estimation of neighborhoods from descriptors, further highlighting the discussed vulnerability of obfuscation schemes. Future methods in de-noising the neighborhood graphs estimated from descriptor and/or geometry can help reduce the error in point position recovery. Similarly, more sophisticated inversion attacks that are robust to small noise in point positions can increase the privacy risk.
@@ 7 Conclusion
In this work, we highlight a common vulnerability of all geometry-based obfuscation techniques that have so far been presented as privacy-preserving representations. We present a simple optimization-based method that uses knowledge of point neighborhoods to recover point positions from the discussed obfuscation schemes. We show the robustness of our method and analyze the recovery accuracy by using oracle-provided neighborhoods with varying inlier ratios. Finally, using a neural network that learns to identify local feature descriptors co-occurring across scenes, we show that it is possible to estimate these neighborhoods from the descriptors associated with points. The inverted images from the recovered point positions reveal private scene content, highlighting the drawback of current methods and the need for guarantees on under which circumstances a data representation is indeed privacy-preserving.
Acknowledgements.  This work was supported by the Czech Science Foundation (GACR) JUNIOR STAR Grant No. 22-23183M, the EU Horizon 2020 project RICAIP (No. 857306), the Chalmers AI Research Center (CHAIR), WASP and SSF. The compute and storage were partially supported by NAISS projects numbered NAISS-2024/22-637 and NAISS-2024/22-237 respectively.
[…]
@@
The supplementary material is organized as follows. Sec. A details the point recovery from coordinate permutations [44] and how we estimate which of the coordinates is swapped to transform the recovery from coordinate permutations into a recovery from lines. Sec. B recalls the descriptor ambiguity in paired-point lines obfuscations [35] and how neighborhood information is used to assign descriptors to their original points. Sec. C reports results on the indoor 12-scenes [68] dataset as announced in Section 6. These results are consistent with the ones on the indoor 7-scenes [60] dataset. We also report the geometric and perceptual evaluation for all 3D obfuscations, including the random line obfuscation OLC [62] and the PPL+ variant of the pair-point lifting [35], that are left out of the main paper for the sake of brevity. Additionally, we also provide visual examples of the estimated neighborhood graph on two scenes from the ScanNet++ [73] dataset. Sec. D provides additional implementation details related to the nearest-neighbor learning and the image inversion from 2D points.
Appendix A Coordinate permutation - Predicting swapped coordinate
@@
Appendix B Descriptor Assignment for Paired-Point-Lines
@@ Appendix D Implementation Details
For 3D lines [62, 35, 42] and 3D lines made from 3D permutation [44], the 3D points to recover are initialized as the projection of a 3D point ”anchor” onto the lines. The 3D anchor is defined as follows: the 3D lines are projected onto a plane. We sample a set of intersections between the resulting 2D lines and compute their 2D centroid, which we use as the anchor. In the paper, we use the plane $z=0$ so the centroid has the form $(x,y,0)$ and randomly sample 10K intersection points. Note that the choice for the plane $z=0$ does not necessarily correspond to the ground-plane as the coordinate frames of the scenes are chosen arbitrarily by the authors of the datasets.
@@
For planes, we also project a 3D ”anchor” point onto each plane but the anchor point is built differently: it is defined as the 3D point which coordinates are the average of the planes’ offsets associated with that axis, i.e., the $x$ coordinate is the average of offsets $c$ of all planes of the form $x=c$.
@@
Visualizing Estimated Neighbors.  As an additional way to to evaluate the quality of the estimated neighborhood, we draw the neighborhood graph on top of the images in Figures 17 and  18. SuperPoint [16] keypoints form the nodes of the graph and the graph has an edge between each point and its top-5 nearest neighbors estimated by our neighborhood estimation network (Sec. 5 of the main paper). We use two scenes taken from the ScanNet++ [73] dataset showing a bedroom and an office. The accuracy of our proposed point position estimation depends on the distance of the nearest neighbors used – the error in estimation increases if points that are far apart are considered as neighbors. We therefore color the edges green if the distance between the corresponding nodes is less than a threshold and red otherwise. We use $\epsilon=0.1*min(h,w)$ as the threshold where $h,w$ are the height and width of the image. It is worth noting that in regions of images with more texture, such as texts, paintings, and other distinct objects, the keypoint density is high and the quality of estimated neighbors is also high. These are regions typically containing private user content. In texture-less parts of the scene, such as floors, walls and ceilings, the keypoints are sparse and their local regions are visually less distinct, making neighborhood estimation difficult, as illustrated by several red edges. However, often such regions do not contain information that is private to the user. Measures such as SSIM and PSNR treat all parts of the image equally, whereas from a privacy point of view, recovering certain parts of the image with good detail is enough to deem the method as not privacy-preserving. More nuanced methods to measure the privacy aspect of inverted images are therefore needed.
Figure 17: Neighborhood estimated from SuperPoint  [16] descriptors: Images from scene 0a76e0647 from the ScanNet++ [73] showing detected SuperPoint  [16] keypoints and the neighborhood graph estimated using our network described in Sec. 5 of the main paper. Top-5 neighbors for each point have been plotted with edges colored green if the points are closer than a threshold and red otherwise.
Figure 18: Neighborhood estimated from SuperPoint  [16] descriptors: Images from scene 036bce3393 from the ScanNet++ [73] showing detected SuperPoint  [16] keypoints and the neighborhood graph estimated using our network described in Sec. 5 of the main paper. Top-5 neighbors for each point have been plotted with edges colored green if the points are closer than a threshold and red otherwise.